how do i work this problem 1/4(x-3)+5/6=7/12x+2/3

i am stuck, the lcd is 12 so do i divide 12 on both sides?

Answer by Alan3354(69443)

how do i work this problem 1/4(x-3)+5/6=7/12x+2/3
i am stuck, the lcd is 12 so do i divide 12 on both sides?
----------------
It's not clear if it's 1/(4(x+3)) or (1/4)(x+3), but since you say the LCD is 12:
(x+3)/4 + 5/6 = 7x/12 + 2/3
Multiply by 12
3(x+3) + 10 = 7x + 8
3x+9 + 10 = 7x+8
3x+19 = 7x+8
4x = 11
x = 11/4